# HyPINO: Multi-Physics Neural Operators via HyperPINNs and the Method of Manufactured Solutions

### v3

@@ 4.2 Baseline Models
We compare HyPINO against three baselines, each trained for 30,000 batches with batch size 128 and an initial learning rate of $10^{-4}$: (i) UNet [42], which shares HyPINO’s encoder but replaces the hypernetwork decoder with a convolutional decoder that directly outputs a $224\times 224$ solution grid. It is trained solely on supervised data and has 62M trainable parameters. (ii) Poseidon [19], a pretrained neural operator with 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the embedding and lead-time-conditioned layer normalization layers’ dimensionality to match the size of our parametrization. Similarly to the UNet, Poseidon is fine-tuned only on supervised data. (iii) PINO [29], a Fourier neural operator [28] with 33M parameters. We adapt it to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It is trained using the same hybrid supervision and curriculum as HyPINO, including physics-informed losses.
@@ 4.3 Evaluation
We evaluate HyPINO and baseline models on seven standard PDE benchmarks from the PINN literature: (i) HT - 1D heat equation [32], (ii) HZ - 2D Helmholtz equation [2], (iii) HZ-G - Helmholtz on an irregular geometry [16], (iv) PS-C - Poisson with four circular interior boundaries [16], (v) PS-L - Poisson on an L-shaped domain [32], (vi) PS-G - Poisson with a Gaussian vorticity field [19], and (vii) WV - 1D wave equation [16]. The exact problem statements and visualizations of the corresponding parameterizations are provided in Appendix B.
Table 1: Model performance across seven PDE benchmarks. Each cell shows mean squared error (MSE) / symmetric mean absolute percentage error (SMAPE) [34]. Lower is better.
@@
We summarize model performance across the seven PDE benchmarks in Table 1. HyPINO demonstrates consistently strong results, achieving an average rank of 2.00 across all tasks, compared to 3.00 for UNet, 2.86 for Poseidon, and 2.14 for PINO. It is important to note that neither Poseidon nor PINO was originally designed for the PDE parameterization chosen in this study. As such, some degree of performance degradation is expected. In contrast, HyPINO contains a dedicated embedding mechanism tailored to this parameterization, but faces the challenge of operating in a significantly less structured output space compared to the grid-based outputs of the baselines. Its competitive zero-shot performance under these conditions is therefore noteworthy. Across all benchmarks, models trained with physics-informed objectives generally outperform those relying solely on supervised data. This indicates that incorporating physics-based losses helps mitigate the generalization gap between the synthetic training data and the evaluation tasks.
Table 1 further highlights the advantages of our proposed iterative refinement approach. After three refinement iterations (HyPINO3), we observe substantial reductions in prediction error across all but one benchmark. Notably, the MSE for PS-C and PS-G decreases by more than one order of magnitude, and an even larger improvement is observed for HT (almost two orders of magnitude). With ten refinement iterations (HyPINO10), our model achieves state-of-the-art performance on all but two evaluated benchmarks, outperforming the best baseline models by factors ranging from 2.1 (on HZ against Poseidon) to 173 (HT against PINO). Table 2 shows that iterative refinement leads to a continuously more accurate prediction on the challenging WV benchmark, with the model being able to extend the undulating shape continuously further from the initial condition across the time dimension. Importantly, our results indicate that iterative refinement is not specific to HyPINO but serves as a generally effective test-time enhancement for other physics-informed neural operators, as demonstrated by the performance of PINO3 and PINO10.
We hypothesize that these improvements arise because the iterative procedure allows for correcting systematic biases introduced during training on synthetic data, which, despite its diversity and breadth, remains composed of relatively simple basis functions. As these training-induced errors tend to be consistent, artifacts produced by the initial HyPINO-generated PINNs can be systematically corrected in subsequent iterations. This residual-driven refinement yields ensembles that are significantly more effective than naive ensembles formed from independently trained models.
@@
4.4 Finetuning
The parameters $\theta^{\star}$ produced by HyPINO can be used to initialize PINNs for subsequent fine-tuning on specific PDE instances. We compare the convergence behavior of HyPINO-initialized PINNs with those initialized randomly and with Reptile meta-learning [38], where Reptile was trained on our synthetic dataset using 10,000 outer- and 1,000 inner-loop cycles. We also evaluate ensembles generated with HyPINO3 and HyPINO10 against ensembles of equal size and architecture initialized with random weights or Reptile.
@@ A.1 Neural Operator Architecture
We build on the HyperPINN framework [8] and design a neural operator based on a hypernetwork that generates the weights $\theta^{\star}$ of a Physics-Informed Neural Network (PINN) $u_{\theta}$, conditioned on a given PDE instance. Specifically, the hypernetwork learns a mapping
@@
Each grid-valued input is first passed through a Fourier feature mapping [45], which augments the input with sinusoidal encodings using five exponentially increasing bands $\text{frequencies}=0.1\cdot 2^{i},\quad i\in\{0,1,2,3,4\}$. This enhances the network’s ability to represent high-frequency content and reduces spectral bias. The Fourier mapping layer is followed by two convolutional layers with kernel size three and strides of two. For the boundary location grids $M$ (c.f. Section 3.1, we compute four embeddings: $z_{D}^{1},\;z_{D}^{2},\;z_{N}^{1},\;z_{N}^{2}$. For the boundary value grids $V$, we compute $z_{g}$ (Dirichlet values) and $z_{h}$ (Neumann values). The source term yields the embedding $z_{f}$.
@@
The grid embedding $z_{G}$ is processed by a sequence of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$. Denoting by $z^{(i)}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ the output of block $\mathcal{SW}_{i}$ and $z^{(0)}=z_{G}$, we interleave each block with a FiLM modulation [40] conditioned on the coefficient embeddings $z_{C}$. Concretely, we define
@@
Inspired by Swin Transformer UNet architectures [4, 11], we retain all intermediate latent representations from the Swin blocks $\{z^{(i)}\}_{i=1}^{K}$ to keep information at various semantic levels.
@@
To aggregate spatial information into a compact latent representation suitable for parameterizing the target PINN, we perform Multi-Head Attention Pooling [25, 55] across the flattened outputs of each Swin Transformer block. Specifically, let $z_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ denote the output of the $i$-th FiLM-modulated Swin block. We reshape it into a sequence of tokens $kv_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$, which serve as both keys and values in the attention mechanism.
@@
We define the architecture of the target PINN as MLP with Fourier feature mapping [45] and multiplicative skip connections [46]. Fourier encodings provide spectral expressivity for modeling high-frequency components [47], while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by masking some layers.
@@
Following Wang et al. [46], the encoded input is projected through three parallel transformations:
@@ A.2.2 Sampling Differential Operators
Considering $\mathcal{B}=\{u,u_{x},u_{y},u_{xx},u_{yy}\}$ to be the set of all terms that can appear in our differential operators, we sample the number of terms $n\sim\text{Uniform}({1,2,3})$. We then randomly select $n$ terms from $\mathcal{B}$ without repetition and obtain their coefficients (c.f. Section 3.1) $c_{i}\sim\text{Uniform}([-2,2])$. The sum of the selected terms multiplied by their respective coefficients constitutes the final differential operator.
@@ A.2.5 Sampling Physical Domains
We employ a randomized sampling procedure based on Constructive Solid Geometry (CSG) [33] to generate complex and diverse domains.
@@ A.2.6 Sampling Boundary Conditions
For inner boundaries (created by subtracting geometric shapes via CSG, c.f. Section A.2.5), each component $\partial\Omega_{\text{inner},i}$, with $i\in\{0,1,\dots,n\}$, is independently assigned either a Dirichlet or Neumann condition, or both:
@@ B.1 Heat 1D (HT)
This benchmark problem is adapted from DeepXDE [32]. Figure 5 shows the parametrization of the different PDE components.
@@
Figure 5: Parametrization of the 1D Heat PDE.
@@ B.2 Helmholtz 2D (HZ)
This benchmark problem is adapted from DeepXDE [32]. Figure 6 shows the parametrization of the different PDE components.
@@
Figure 6: Parametrization of the 2D Helmholtz PDE.
@@ B.3 Helmholtz 2D - Complex Geometry (HZ-G)
This benchmark problem is adapted from PINNacle [16]. Figure 7 shows the parametrization of the different PDE components.
@@
Figure 7: Parametrization of the 2D Helmholtz PDE with complex geometry.
@@ B.4 Poisson 2D - Circles (PS-C)
This benchmark problem is adapted from PINNacle [16]. Figure 8 shows the parametrization of the different PDE components.
@@
Figure 8: Parametrization of the 2D Poisson PDE with circular inner boundaries.
@@ B.5 Poisson 2D - L-Domain (PS-L)
This benchmark problem is adapted from DeepXDE [32]. Figure 9 shows the parametrization of the different PDE components.
@@
Figure 9: Parametrization of the 2D Poisson PDE on an L-shaped domain.
@@ B.6 Poisson 2D - Gauss (PS-G)
We select a sample from the dataset introduced in [19]. Figure 10 shows the parametrization of the different PDE components.
@@
Figure 10: Parametrization of the 2D Poisson PDE with Gaussian superposition vorticity field.
@@ B.7 Wave 1D (WV)
This benchmark problem is adapted from PINNacle [16]. Figure 11 shows the parametrization of the different PDE components.
@@
Figure 11: Parametrization of the 1D Wave PDE.
@@ B.8 Baseline Models
UNet [42]. A convolutional encoder–decoder network that shares the same input encoding architecture as HyPINO, but replaces the transformer-based decoder with a purely convolutional upsampling stack that directly outputs a solution grid of shape $(224\times 224)$, matching the resolution of the input tensors. It is trained exclusively on supervised PDEs with analytical solutions, using a batch size of 128, an initial learning rate of $10^{-4}$, and for 30,000 training batches. The UNet has a total parameter count of 62M.
Poseidon [19]. A large pretrained operator network with approximately 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the input dimensionality to 5 to accept all grid-based inputs. Additionally, the lead-time-conditioned layer normalization layers—originally designed to condition on a 1D time input—are modified to condition on the 5D vector of differential operator coefficients. Poseidon is fine-tuned exclusively on supervised samples, using the same training setup as the UNet (30,000 batches, batch size 128, initial learning rate $10^{-4}$).
PINO[29]. A Fourier neural operator[28] architecture with 33M parameters, trained with joint physics-informed and supervised losses computed in Fourier space. We adapt the model to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It follows the same hybrid supervision and training curriculum as HyPINO, including physics-informed losses, and is trained for 30,000 batches with a batch size of 128 and an initial learning rate of $10^{-4}$.
@@ B.9 Evaluation
B.10 Finetuning
We investigate the utility of HyPINO-generated PINN parameters $\theta^{\star}$ as a prior for rapid adaptation to specific PDE instances. We compare three initialization strategies: (i) HyPINO-initialized PINNs, (ii) randomly initialized PINNs, and (iii) PINNs initialized via Reptile meta-learning [38]. For Reptile, we use 10,000 outer-loop steps on our synthetic dataset and 1,000 inner-loop updates per sample with an inner learning rate of 0.01.
In addition to single-network performance, we evaluate the effect of initialization on ensemble-based methods. We compare ensembles of size 3 and 10 generated using HyPINO (denoted HyPINO3 and HyPINO10) against ensembles of the same size and architecture initialized either randomly or via Reptile. For the latter, we replicate the Reptile-initialized weights across all ensemble members. Note that a HyPINOi ensemble consists of the base PINN as well as three refinement (or delta) PINNs: $u^{(0)}+\sum_{t=1}^{T}\delta u^{(t)}$, thus creating an ensemble with four experts.

### v4

@@ 4.2 Baseline Models
We compare HyPINO against three baselines, each trained for 30,000 batches with batch size 128 and an initial learning rate of $10^{-4}$: (i) U-Net [unet], which shares HyPINO’s encoder but replaces the hypernetwork decoder with a convolutional decoder that directly outputs a $224\times 224$ solution grid. It is trained solely on supervised data and has 62M trainable parameters. (ii) Poseidon [pde_foundation_model_2024_Herde], a pretrained neural operator with 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the embedding and lead-time-conditioned layer normalization layers’ dimensionality to match the size of our parameterization. Similarly to the U-Net, Poseidon is fine-tuned only on supervised data. (iii) PINO [pino], a Fourier neural operator [fno] with 33M parameters. We adapt it to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It is trained using the same hybrid supervision and curriculum as HyPINO, including physics-informed losses.
@@ 4.3 Evaluation
We evaluate HyPINO and baseline models on seven standard PDE benchmarks from the PINN literature: (i) HT - 1D heat equation [deepxde], (ii) HZ - 2D Helmholtz equation [relobralo], (iii) HZ-G - Helmholtz on an irregular geometry [pinnacle], (iv) PS-C - Poisson with four circular interior boundaries [pinnacle], (v) PS-L - Poisson on an L-shaped domain [deepxde], (vi) PS-G - Poisson with a Gaussian vorticity field [pde_foundation_model_2024_Herde], and (vii) WV - 1D wave equation [pinnacle]. The exact problem statements and visualizations of the corresponding parameterizations are provided in Appendix B.
Table 1: Model performance across seven PDE benchmarks. Each cell shows mean squared error (MSE) / symmetric mean absolute percentage error (SMAPE) [smape]. Lower is better.
@@
We summarize model performance across the seven PDE benchmarks in Table 1. HyPINO demonstrates consistently strong results, achieving an average rank of 2.00 across all tasks, compared to 3.00 for U-Net, 2.86 for Poseidon, and 2.14 for PINO. It is important to note that neither Poseidon nor PINO was originally designed for the PDE parameterization chosen in this study. As such, some degree of performance degradation is expected. In contrast, HyPINO contains a dedicated embedding mechanism tailored to this parameterization, but faces the challenge of operating in a significantly less structured output space compared to the grid-based outputs of the baselines. Its competitive zero-shot performance under these conditions is therefore noteworthy. Across all benchmarks, models trained with physics-informed objectives generally outperform those relying solely on supervised data. This indicates that incorporating physics-based losses helps mitigate the generalization gap between the synthetic training data and the evaluation tasks.
Table 1 further highlights the advantages of our proposed iterative refinement approach. After three refinement iterations (HyPINO3), we observe substantial reductions in prediction error across all but one benchmark. Notably, the MSE for PS-C and PS-G decreases by more than one order of magnitude, and an even larger improvement is observed for HT (almost two orders of magnitude). With ten refinement iterations (HyPINO10), our model achieves state-of-the-art performance on all but two evaluated benchmarks, outperforming the best baseline models by factors ranging from 2.1 (on HZ against Poseidon) to 173 (HT against PINO). Table 2 shows that iterative refinement leads to a progressively more accurate prediction on the challenging WV benchmark, with the model being able to extend the undulating shape continuously further from the initial condition across the time dimension. Importantly, our results indicate that iterative refinement is not specific to HyPINO but serves as a generally effective test-time enhancement for other physics-informed neural operators, as demonstrated by the performance of PINO3 and PINO10.
We hypothesize that these improvements arise because the iterative procedure allows for correcting systematic biases introduced during training on synthetic data, which, despite its diversity and breadth, remains composed of relatively simple basis functions. As these training-induced errors tend to be consistent, artifacts produced by the initial HyPINO-generated PINNs can be systematically corrected in subsequent iterations. This residual-driven refinement yields ensembles that are significantly more effective than naive ensembles formed from independently generated target networks.
@@
4.4 Fine-tuning
The parameters $\theta^{\star}$ produced by HyPINO can be used to initialize PINNs for subsequent fine-tuning on specific PDE instances. We compare the convergence behavior of HyPINO-initialized PINNs with those initialized randomly and with Reptile meta-learning [reptile], where Reptile was trained on our synthetic dataset using 10,000 outer- and 1,000 inner-loop cycles. We also evaluate ensembles generated with HyPINO3 and HyPINO10 against ensembles of equal size and architecture initialized with random weights or Reptile.
@@ A.1 Neural Operator Architecture
We build on the HyperPINN framework [hyperpinn] and design a neural operator based on a hypernetwork that generates the weights $\theta^{\star}$ of a Physics-Informed Neural Network (PINN) $u_{\theta}$, conditioned on a given PDE instance. Specifically, the hypernetwork learns a mapping
@@
Each grid-valued input is first passed through a Fourier feature mapping [tancik2020fourier], which augments the input with sinusoidal encodings using five exponentially increasing bands $\text{frequencies}=0.1\cdot 2^{i},\quad i\in\{0,1,2,3,4\}$. This enhances the network’s ability to represent high-frequency content and reduces spectral bias. The Fourier mapping layer is followed by two convolutional layers with kernel size three and strides of two. For the boundary location grids $M$ (cf. Section 3.1), we compute four embeddings: $z_{D}^{1},\;z_{D}^{2},\;z_{N}^{1},\;z_{N}^{2}$. For the boundary value grids $V$, we compute $z_{g}$ (Dirichlet values) and $z_{h}$ (Neumann values). The source term yields the embedding $z_{f}$.
@@
The grid embedding $z_{G}$ is processed by a sequence of $K$ Swin Transformer blocks $\{\mathcal{SW}_{i}\}_{i=1}^{K}$. Denoting by $z^{(i)}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ the output of block $\mathcal{SW}_{i}$ and $z^{(0)}=z_{G}$, we interleave each block with a FiLM modulation [film] conditioned on the coefficient embeddings $z_{C}$. Concretely, we define
@@
Inspired by Swin Transformer U-Net architectures [swin_unet, swin_unet_denoising], we retain all intermediate latent representations from the Swin blocks $\{z^{(i)}\}_{i=1}^{K}$ to keep information at various semantic levels.
@@
To aggregate spatial information into a compact latent representation suitable for parameterizing the target PINN, we perform Multi-Head Attention Pooling [lee2019set, zhai2022scaling] across the flattened outputs of each Swin Transformer block. Specifically, let $z_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ denote the output of the $i$-th FiLM-modulated Swin block. We reshape it into a sequence of tokens $kv_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$, which serve as both keys and values in the attention mechanism.
@@
We define the architecture of the target PINN as an MLP with Fourier feature mapping [tancik2020fourier] and multiplicative skip connections [wang2021mult_skip_connections]. Fourier encodings provide spectral expressivity for modeling high-frequency components [fourier_pinns], while the skip connections enhance gradient propagation and, in the context of hypernetworks, have the additional benefit of enabling dynamic depth modulation based on PDE complexity by masking some layers.
@@
Following wang2021mult_skip_connections, the encoded input is projected through three parallel transformations:
@@ A.2.2 Sampling Differential Operators
Considering $\mathcal{B}=\{u,u_{x},u_{y},u_{xx},u_{yy}\}$ to be the set of all terms that can appear in our differential operators, we sample the number of terms $n\sim\text{Uniform}({1,2,3})$. We then randomly select $n$ terms from $\mathcal{B}$ without repetition and obtain their coefficients (cf. Section 3.1) $c_{i}\sim\text{Uniform}([-2,2])$. The sum of the selected terms multiplied by their respective coefficients constitutes the final differential operator.
@@ A.2.5 Sampling Physical Domains
We employ a randomized sampling procedure based on Constructive Solid Geometry (CSG) [deepxde] to generate complex and diverse domains.
@@ A.2.6 Sampling Boundary Conditions
For inner boundaries (created by subtracting geometric shapes via CSG, cf. Section A.2.5), each component $\partial\Omega_{\text{inner},i}$, with $i\in\{0,1,\dots,n\}$, is independently assigned either a Dirichlet or Neumann condition, or both:
@@ B.1 Heat 1D (HT)
This benchmark problem is adapted from DeepXDE [deepxde]. Figure 5 shows the parameterization of the different PDE components.
@@
Figure 5: Parameterization of the 1D Heat PDE.
@@ B.2 Helmholtz 2D (HZ)
This benchmark problem is adapted from DeepXDE [deepxde]. Figure 6 shows the parameterization of the different PDE components.
@@
Figure 6: Parameterization of the 2D Helmholtz PDE.
@@ B.3 Helmholtz 2D - Complex Geometry (HZ-G)
This benchmark problem is adapted from PINNacle [pinnacle]. Figure 7 shows the parameterization of the different PDE components.
@@
Figure 7: Parameterization of the 2D Helmholtz-type (Poisson–Boltzmann) PDE with complex geometry.
@@ B.4 Poisson 2D - Circles (PS-C)
This benchmark problem is adapted from PINNacle [pinnacle]. Figure 8 shows the parameterization of the different PDE components.
@@
Figure 8: Parameterization of the 2D Poisson PDE with circular inner boundaries.
@@ B.5 Poisson 2D - L-Domain (PS-L)
This benchmark problem is adapted from DeepXDE [deepxde]. Figure 9 shows the parameterization of the different PDE components.
@@
Figure 9: Parameterization of the 2D Poisson PDE on an L-shaped domain.
@@ B.6 Poisson 2D - Gauss (PS-G)
We select a sample from the dataset introduced in [pde_foundation_model_2024_Herde]. Figure 10 shows the parameterization of the different PDE components.
@@
Figure 10: Parameterization of the 2D Poisson PDE with Gaussian superposition vorticity field.
@@ B.7 Wave 1D (WV)
This benchmark problem is adapted from PINNacle [pinnacle]. Figure 11 shows the parameterization of the different PDE components.
@@
Figure 11: Parameterization of the 1D Wave PDE.
@@ B.8 Baseline Models
U-Net [unet]. A convolutional encoder–decoder network that shares the same input encoding architecture as HyPINO, but replaces the transformer-based decoder with a purely convolutional upsampling stack that directly outputs a solution grid of shape $(224\times 224)$, matching the resolution of the input tensors. It is trained exclusively on supervised PDEs with analytical solutions, using a batch size of 128, an initial learning rate of $10^{-4}$, and for 30,000 training batches. The U-Net has a total parameter count of 62M.
Poseidon [pde_foundation_model_2024_Herde]. A large pretrained operator network with approximately 158M parameters. We use the Poseidon-B checkpoint and adapt it by changing the input dimensionality to 5 to accept all grid-based inputs. Additionally, the lead-time-conditioned layer normalization layers—originally designed to condition on a 1D time input—are modified to condition on the 5D vector of differential operator coefficients. Poseidon is fine-tuned exclusively on supervised samples, using the same training setup as the U-Net (30,000 batches, batch size 128, initial learning rate $10^{-4}$).
PINO [pino]. A Fourier neural operator [fno] architecture with 33M parameters, trained with joint physics-informed and supervised losses computed in Fourier space. We adapt the model to accept 5-channel grid inputs and condition on the PDE operator using FiLM layers. It follows the same hybrid supervision and training curriculum as HyPINO, including physics-informed losses, and is trained for 30,000 batches with a batch size of 128 and an initial learning rate of $10^{-4}$.
@@ B.9 Evaluation
B.10 Fine-tuning
We investigate the utility of HyPINO-generated PINN parameters $\theta^{\star}$ as a prior for rapid adaptation to specific PDE instances. We compare three initialization strategies: (i) HyPINO-initialized PINNs, (ii) randomly initialized PINNs, and (iii) PINNs initialized via Reptile meta-learning [reptile]. For Reptile, we use 10,000 outer-loop steps on our synthetic dataset and 1,000 inner-loop updates per sample with an inner learning rate of 0.01.
In addition to single-network performance, we evaluate the effect of initialization on ensemble-based methods. We compare ensembles of size 3 and 10 generated using HyPINO (denoted HyPINO3 and HyPINO10) against ensembles of the same size and architecture initialized either randomly or via Reptile. For the latter, we replicate the Reptile-initialized weights across all ensemble members. Note that a HyPINOi ensemble consists of the base PINN as well as $i$ refinement (or delta) PINNs: $u^{(0)}+\sum_{t=1}^{T}\delta u^{(t)}$, thus creating an ensemble with $i+1$ experts.